OpenStudy (anonymous):

Solve the inequality. -2/5x - 9 < 9/10

OpenStudy (cloverracer):

What are the answer choices?

OpenStudy (anonymous):

there is none.

OpenStudy (cloverracer):

Okay, do you think you can enter this in equation form? (The blue button below that says Equation)

OpenStudy (anonymous):

\[-2/5x - 9 < 9/10 \]

OpenStudy (anonymous):

\[-\frac{ 2 }{ 5 }x - 9 < \frac{ 9 }{ 10 }\]

OpenStudy (anonymous):

is that better?

OpenStudy (cloverracer):

Yes, thanks! I will solve it.

OpenStudy (anonymous):

okay thanks

OpenStudy (cloverracer):

Okay, I got. \[x < \frac{ 4 }{ 99 } \] or \[x > 0\]

OpenStudy (anonymous):

Okay thanks!!

OpenStudy (cloverracer):

You're welcome!